OpenStudy (anonymous):

solve the system by substitution 2x+y=-11 3x-4y=11 choices A.(3,5) B.(-5,-3) C.(-3,-5) D.(5,3)

OpenStudy (anonymous):

C. try substituting the values in option C

OpenStudy (anonymous):

okay what about solve the system using elimination 2x+6y=-12 5x-5y=10 choices A.(2,1) B.(0,-2) C.(-2,0) D.(1,2)

OpenStudy (anonymous):

For elimination methosd, 2x+6y=-12 ---> (1) 5x-5y=10 --->(2) Let's make both eqn (1) and eqn (2) have 10x in them. to do so, multiply. So for (1), multiply by 5. it'll become 10x+30y=-60 ---> (3) for (2), multiply by 2. it'll become 10x-10y=20 --->(4)

OpenStudy (anonymous):

Then to eliminate 10x, take (3)-(4) so, 10x+30y-(10x-10y) = -60 - 20 40y=-80 y=-2 :) then sub y=-2 into (1) 2x+6(-2)=-12 2x=0 x=0 :)

OpenStudy (anonymous):

do u understand?

OpenStudy (anonymous):

okay so you jus make both equations have a common number in them

OpenStudy (anonymous):

yes that's right

OpenStudy (anonymous):

okay thanks!!

OpenStudy (anonymous):

np

OpenStudy (zpupster):

i was watching , i would like to comment on the first problem "solve using substitution.

OpenStudy (anonymous):

what about it?

OpenStudy (zpupster):

you do not just substitute given answers to see if it satisfies the equation you do it like this 2x+y=-11 3x-4y=11 first solve for a variable , y would be easiest because it is by itself in equation 1 y= -11 -2x subtracted 2x from both sides and then substitute that in for y in the 2nd equation like so: 3x- 4( -11 -2x) =11 and then solve

OpenStudy (zpupster):

once you get an answer for x plug that back in to get y and do not forget to do your checks

OpenStudy (anonymous):

do you mulitply -4 and -11 and -4 and -2x?

OpenStudy (zpupster):

yes

OpenStudy (anonymous):

okay so you get 3x-44-8x=11 do you combine like terms (the 3x and -8x)

OpenStudy (zpupster):

-4 times -2x is positive

OpenStudy (zpupster):

3x and 8x pay careful attention to your signs

OpenStudy (anonymous):

so now i have 11x+44=11 subtract 44 from both sides and get 11x=-33

OpenStudy (zpupster):

yes

OpenStudy (anonymous):

now what

OpenStudy (zpupster):

divide both sides by 11 to isolate x

OpenStudy (anonymous):

x=-3

OpenStudy (anonymous):

so the anwser is C

OpenStudy (zpupster):

yes take that and plug it in on of the equations and solve for y

OpenStudy (anonymous):

alright thanks for your help

OpenStudy (zpupster):

yes good job!!!! remember to plug your x and y values in both equations as a check